OpenStudy (kweenlexis):

A recipe calls for 4 L of water. Arlene wants to make 0.3 more servings than the recipe makes. Answer in comments. Which expression represents how many liters of water Arlene should use?

OpenStudy (denonakavro):

shouldnt it be (4 x 0.3)+4

OpenStudy (kweenlexis):

A. \[\frac{ 3 }{ 10} \times4\] B. \[\frac{ 13 }{ 10} \times 4\] C. \[\frac{ 7 }{ 10 } \times 4\] D. \[\frac{ 17 }{ 10 }\times4\]

OpenStudy (trantom):

it depends on how many servers 4 L gives you

OpenStudy (trantom):

*servings

OpenStudy (denonakavro):

it is b

OpenStudy (trantom):

well for that it's b

OpenStudy (kweenlexis):

ok thanks

OpenStudy (denonakavro):

np